OpenStudy (anonymous):

The slope-intercept form of the equation of a line that passes through point (–3, 8) is y = -2/3x + 6. What is the point-slope form of the equation for this line? y – 3 = –2/3(x + 8) y + 3 = –2/3(x – 8) y + 8 = –2/3(x – 3) y – 8 = –2/3(x + 3)

OpenStudy (texaschic101):

point slope form : y - y1 = m(x - x1) y = - 2/3x + 6...slope(m) is -2/3 (-3,8)...x1 = -3 and y1 = 8 now we sub y - 8 = -2/3(x - (-3) = y - 8 = -2/3(x + 3)

OpenStudy (anonymous):

so its d

OpenStudy (anonymous):

The point-slope form of the equation of the line that passes through (–5, –1) and (10, –7) is y + 7 = (x – 10). What is the standard form of the equation for this line? 2x – 5y = –15 2x – 5y = –17 2x + 5y = –15 2x + 5y = –17